OpenStudy (anonymous):

Help with maths? If f(x) = 4x - 3 and g(x) = 2x + 2, find the value of: a. f(-1) b. g(0) c. f(2) d. f^-1(x) e. fg(x) f. gf(x) g. If f(x) = 5, find x

OpenStudy (anonymous):

ok lets break it down. Are you ready?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

wait, are A-G possible answers or are those the questions?

OpenStudy (anonymous):

They are the questions I have to solve

OpenStudy (anonymous):

ok Lets do one. Pick with one you want me to help you with.

OpenStudy (anonymous):

E

OpenStudy (anonymous):

@mathmale Let me get his help he is better at this then me.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Lets do A first

OpenStudy (anonymous):

Okay sure

OpenStudy (anonymous):

oh...im sorry I can't do thease for you I don't understand it, please make a new question so others can help you :( sorry again :'(

OpenStudy (anonymous):

No problem

OpenStudy (gorv):

okkk listen

OpenStudy (gorv):

f(-1) means we put x=-1 in f(x) so in the function replace x by -1

OpenStudy (gorv):

for f(-1) f(-1)=4*(-1)-3=-4-3=-7 similarly do b and c thn i will tell the next

OpenStudy (anonymous):

I've done b and c

OpenStudy (anonymous):

d, e and f is my problem